A rectangular room has a perimeter of 48ft. The width is 2/3 of the length. What are the dimensions of the room

Answer by dileepkumarabhijith(12)

hi

perimeter 2*(l+b)=48
l+b= 48/2=24
l+2/3l=24
5l/3=24
5l=72
l=72/5=14.4

b=2/3*l=2/3*14.4=9.6